1. 200 mL of an aqueous solution contains 0.030 M concentrations of both Pb2+ and Ag+. If 100 mL of 6.0 x 10-2 M NaCl is added to this solution will a precipitate form? If so, what will the precipitate be? The Ksp values for PbCl2 and AgCl are 1.7 x 10-5 and 1.8 x 10-
Answer #1

No. of moles of Pb2+ = Molarity x Volume

= 0.03 M x 2 L = 0.06 mol

No. of moles of Ag2+ = Molarity x Volume

= 0.03 M x 2 L = 0.06 mol

No. of moles of Cl- = Molarity x Volume

= 0.06 M x 1 L = 0.06 mol

Total volume of solution = 200 mL + 100 mL = 300 mL = 3 L

[Pb2+] = No. of moles / Volume (L) = 0.06 mol / 3 L = 0.02 M

[Ag2+] = No. of moles / Volume (L) = 0.06 mol / 3 L = 0.02 M

[Cl-] = No. of moles / Volume (L) = 0.06 mol / 3 L = 0.02 M

PbCl2 (aq) \rightleftharpoons Pb2+ (aq) + 2 Cl- (aq)

AgCl (aq) \rightleftharpoons Ag+ (aq) + Cl- (aq)

Reaction quotient = QPbCl2 = [Pb2+][Cl-]2

Putting the values we get, QPbCl2 = (0.02)(0.02)2 = 8 x 10-6

Ksp of PbCl2 = 1.7 x 10-5 = 17 x 10-6

Since Ksp>Q, hence the precipitate PbCl2 will not form.

Reaction quotient = QAgCl = [Ag+][Cl-]

Putting the values we get, QAgCl = (0.02)(0.02)= 4 x 10-4

Ksp of AgCl= 1.8 x 10-10

Since Q>Ksp, hence the precipitate AgClwill form.